OpenStudy (anonymous):

A business has $11,080 to spend on new laptops and tablet computers for its salespeople. The laptops cost $515 each. The tablets cost $285 each. The business wants each salesperson to have either a laptop or a tablet. There are 30 salespeople. How many of each type of computer should the business buy? A. 9 laptops and 21 tablets B. 11 laptops and 19 tablets C. 19 laptops and 11 tablets D. 21 laptops and 9 tablets

OpenStudy (mathstudent55):

Find the cost of each option, and see which option the company can afford.

OpenStudy (anonymous):

it is a 9 laptops 21 tablets

OpenStudy (anonymous):

so the answer is A.

OpenStudy (anonymous):

thnx :)

OpenStudy (anonymous):

welcome

OpenStudy (mathstudent55):

@Napolions How about explaining how you got to that answer?

OpenStudy (anonymous):

well i multi plied 9 times 515 and 21 times 285 and added the answer 4635 and 5985 and got 10620 so it was A.

OpenStudy (mathstudent55):

Your work is correct, but incomplete, and your answer is wrong.

OpenStudy (mathstudent55):

You should try all options before coming to a conclusion.

OpenStudy (anonymous):

how.

OpenStudy (mathstudent55):

A. 9 laptops and 21 tablets = 9 * 515 + 21 * 85 = 10,620 B. 11 laptops and 19 tablets = 11 * 515 + 19 * 285 = 11,080 C. 19 laptops and 11 tablets = 19 * 515 + 11 * 285 = 12,920 D. 21 laptops and 9 tablets = 21 * 515 + 9 * 285 = 13,380

OpenStudy (anonymous):

so was i right

OpenStudy (mathstudent55):

Notice above that options C and D are over the budget, so they are definitely not acceptable.

OpenStudy (mathstudent55):

Options A and B are within budget. Option B is exactly on budget, and option A is a little bit under budget. Since the company has budgeted $11,080 for the computer equipment, and option B is exactly that amount, I think they should use option B.

OpenStudy (anonymous):

ok but this is lower than the buget and will help save money

OpenStudy (anonymous):

and i know because i already did the question

OpenStudy (mathstudent55):

You can argue that they could save some money, but you can also argue that they want to spend the money they budgeted for this expense.

OpenStudy (anonymous):

And got it right so it is A

OpenStudy (mathstudent55):

@Napolions Are you also doing this work in your math studies?

OpenStudy (anonymous):

yes and no i been finished that equaation and also i going to another part / segment

OpenStudy (mathstudent55):

I have a question for you. When we see questions posted on OS, sometimes it's hard to know exactly how to go about helping the poster because there are different ways of solving a problem, and when we just see a problem without knowing exactly the material the poster is currently learning, we are helping but only in a guessing way. This is my question for you. Is this problem being learned in a section that teaches systems of equations?

OpenStudy (anonymous):

yes and no

OpenStudy (mathstudent55):

We took advantage of the fact that answers are given, and we worked backwards. The way to work this problem is to solve a system of equations. Then we compare what we get with the options, and we choose the correct option. Let x = number of laptops, and let y = number of tablets. x + y = 30 515x + 285y = 11080 We will use the substitution method of solving systems of equations. Solve the first equation for x: x = 30 - y Substitute 30 - y for x in the second equation: 515x + 285y = 11080 515(30 - y) + 285y = 11080 15450 - 515y + 285y = 11080 15450 - 230y = 11080 -230y = -4370 y = 19 Now substitute y = 19 in the first equation: x + y = 30 x + 19 = 30 x = 11 Since x = number of laptops and y = number of tablets, the answer is: 11 laptops and 19 tablets The answer is option B. We are not asked to save the company money. We are asked to find what they can buy. If the idea were to save the company money, the answer would be "buy 30 tablets." Since tablets are only $285 each, 30 tablets would be $8550, and it would save the company even more money. The fact that an option of buying only 30 tablets is not offered shows that saving money is not the goal.